OpenStudy (anonymous):

positive sixteen plus negative thirty-one equals blank negative forty-seven negative fifteen positive fifteen positive forty-seven

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (aggie12341):

negative fifteen

OpenStudy (anonymous):

+16+(-31) =16-31 How to do calculate it? You can calculate 31-16=15 then add a "-" in front of 15. So 16-31=-15 aggie12341 is right.

undeadknight26 (undeadknight26):

can you use a calculator?